Project

General

Profile

Actions

20160927 MIPO: Debrief/actions on MICE Muon Beam

Standing to-do list is in #1876.

Hardware

Target

Target in ISIS running fine, 2.6M actuations. Sudden drop in index level last week, but stable over weekend.

The R78 test stand has been cleaned up; we need to get back to a running state there.

Decay Solenoid

The Kaeser compressor was running hot, hence the air-end replacement in first half of August (to allow the usual recovery and re-work time).
Unable to restart as expected because of the extensive air-con repairs. Eventually diagnosed a leak associated with the infamous bearing; being replaced now.

The two electrical jobs on the BBU circuit remain outstanding; the re-calibration tasks requires the magnet to be ready for powering.
There is a way forward on the Linde PC upgrade but the water débâcle has knocked that back until after the Linde fridge is fixed and stable.

Diffuser

Diffuser seems to have failed around time of power glitch mid-Sept; waiting on tech from Oxford. Failure has two aspects:

  • can't change diffuser settings - not needed (until next week?)
  • broken read-back is not properly ignored by C&M causing Run Control to fail - is a problem now

Conventional Magnets

Regulator board change has got rid of the ramping issues with Q4 - we had a dipole regulator driving a quad. We have a spare quad board in the recent spares purchase and now a spare D1 board. The replacement part for the water-soaked Q2 supply I believe is on order.
The dipole current calibration can be done as long as we have some level of cooling water circulation.

Cooling water

The conventional magnet cooling water has been separated off as a dedicated de-min loop with heat exchanger and pumps in the Trench. The trench circuit itself is mostly ( #1878 ) ready and has demonstrated that it can shift the heat out of the magnets and into the heat exchanger. Significant tasks do remain to be completed (the Earth Leakage measurements are waiting on the conductivity meter, for example) and there are features to add to the PLC.

However we have been unable to keep the magnets stably cool. After raising the Q4 Earth Leakage threshold (this is not a fix!) we have been running with the trench water loop at 30°C on the "cold"/incoming side, and are still limited to 200 MeV settings (the standard reference run settings are marginal on temperature!)

Measurements in eLog 3920 and 3930 show the water in the two circuits leaving the heat exchanger at the same temperature, indicating that we are moving as much heat as we can. Both ends of the exchanger are at a middle temperature, hopefully implying that there's no partial blockage or similar.

Temperature drop off the trench circuit is ~2 °C while the rise in the roof side is ~8-9 °C. Since the flow measured on the trench side is ~250 l/min then the flow through the roof side of the heat exchanger is (estimated) a paltry 60 l/min.

On 19th October 2012 the total flow to conventional magnets was 248 + 84 = 332 l/min.
On 17th June 2015 the total flow to conventional magnets was 209 + 71 = 280 l/min. (quad currents restricted to 315 A)
On 26th June 2016 the total flow to conventional magnets was 148 + 78 = 226 l/min. (quad currents kept restricted to 315 A)
On 29th July 2016 the total flow to conventional magnets was 148 + 75 = 223 l/min. (restricted currents, after re-balancing: eLog 3768 )
On 22nd September 2016 the flow to the conventional magnet heat exchanger was 60 l/min. (quad currents limited around 200 A)

I remarked on the restricted quad currents at CM 43 a year ago - yet by the June data-taking the flows had been reduced further... we managed to recover some of the envelope by re-balancing the water flows, but this won't help now (as the magnet water is all mixed together before going into the heat exchanger). Since then the water was cut back drastically, and then again last week (22nd September 2016) by 3 l/min to the FC cryo-compressors and similar to the dump circuits: 6 l/min is 10% of the flow!

The problem is on the roof side. Plan is to re-plumb the heat exchanger to improve the flow rate to it; if that fails then need to undo split and revert to de-min for entire circuit.

People

BLOCs

Still insufficient volunteers: need five experienced people who can be close to RAL and regularly take 5-7 day stints, and a couple more to cover holes, sickness, etc. Present volunteers inexperienced, distant (limited availability) or have significant other out-of-hours responsibilities. (Been asking since Napa CM, spring 2014). Note the "economic" imperative - people travelling in to stay at RAL would rather get 5 shift credits than 1 BLOC credit. We already have compromised to allow day shifters to cover BLOC, on the understanding that if they are called out overnight then they can take the day following off and be substituted by the stand-by shifter.

Shifters

Shifters continue to need significant training - most of a day and a lot to take in at the best of times. Last week's scheduled, training day was elbowed aside by unprepared Cooling Channel run - became chaotic, disjointed, and training was still not completed when we gave up at 20:30. Wasted Janusz' time, wasted Jaroslaw's time; wasted over a day of my time. Not acceptable and will be raised within University. I don't see how this sort of nonsense helps either shifter or BLOC recruitment.

Updated by Nebrensky, Henry over 6 years ago · 15 revisions